### v1

# Robust and Fine-Grained Detection of AI Generated Texts

Abstract
An ideal detection system for machine generated content is supposed to work well on any generator as many more advanced LLMs come into existence day by day. Existing systems often struggle with accurately identifying AI-generated content over shorter texts. Further, not all texts might be entirely authored by a human or LLM, hence we focused more over partial cases i.e human-LLM co-authored texts. Our paper introduces a set of models built for the task of token classification which are trained on an extensive collection of human-machine co-authored texts, which performed well over texts of unseen domains, unseen generators, texts by non-native speakers and those with adversarial inputs. We also introduce a new dataset of over 2.4M such texts mostly co-authored by several popular proprietary LLMs over 23 languages. We also present findings of our models’ performance over each texts of each domain and generator. Additional findings include comparison of performance against each adversarial method, length of input texts and characteristics of generated texts compared to the original human authored texts.
1 Introduction
Recent advancements in large language models (LLMs) have significantly narrowed the gap between machine-generated and human-authored text. As LLMs continue to improve in fluency and coherence, the challenge of reliably detecting AI-generated content could become increasingly critical. This issue is particularly pressing in domains such as education and online media, where the authenticity of textual material is paramount. While early efforts such as the GLTR (Gehrmann et al., 2019a) provided valuable insights by leveraging statistical methods to differentiate between human and machine text, these methods often lag behind the rapid pace of LLM evolution. Likewise, initiatives aimed at mitigating neural fake news (Zellers et al., 2019a) have made significant strides in addressing the societal implications of AI-generated misinformation. However, as LLMs become more sophisticated, existing detection systems must be re-evaluated and enhanced to maintain their effectiveness. Further, Each domain comes with its version of the issue of detecting machine generated texts. For instance, proprietary LLMs with internet access and better knowledge cutoffs are more likely to be used in domains like academia. Similarly, bad actors might use an open source generators for the task of creating misinformation and deception through machine generated online content as such models can be hosted locally to not leave a trail and are more flexible in terms of not denying user requests. Hence, tailoring models and approaches for each specific domain/scenario might be better applicable for practical scenarios. We chose a token-classification approach to train a model for the task of distinguish writing styles within a text if more than one were found. This approach helped us achieve better performance over texts of unseen features (i.e domain, generator, adversarial inputs, non-native speakers’ texts) as our models were trained to distinguish different styles within a text rather than classifying an input text as one of the two classes it was trained on. Further, we explored the findings and results upon testing our models over other benchmarks which consist of texts from unseen domain and generators. We also tested our models over benchmarks which consist of texts with various adversarial inputs and those written by non-native speakers. We feel our findings and datasets can aid in further research into mitigating the harms of AI generated texts.
2 Related Works
A major portion of current research in detecting machine-generated content focuses on longer-form writing through binary classification. However, AI-generated misinformation is more likely to cause harm than its use in academia, making the distinction between AI and human-generated texts on social media platforms a critical challenge. Existing methods often struggle with accurately identifying AI-generated content over shorter texts. Moreover, binary classification approaches, which categorize texts as either human or AI-generated (Wang et al., 2024a), (Wang et al., 2024b), (Bhattacharjee et al., 2023), (Zellers et al., 2019b), (Macko et al., 2023), (Ghosal et al., 2023), (Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs. In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a), (Ignat et al., 2024), where content typically consists of one or two sentences. Additionally, some detection works rely on detecting watermarks from AI-generated texts, (Chang et al., 2024), (Dathathri et al., 2024), (Sadasivan et al., 2024), (Zhao et al., 2023) but not all generators utilize watermarking limiting the applicability of such approaches. Few other approaches utilize statistical methods (Mitchell et al., 2023), (Kumarage et al., 2023), (Gehrmann et al., 2019b), (Hans et al., 2024), (Bao et al., 2023), but they can be prone to mis-classification against adversarial methods like rephrasing and humanizing. (Abassy et al., 2024) introduced a 4-way classification as entirely human authored, entirely llm authored, human-edited and llm-authored or llm-edited human-authored. An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts. Further an ideal detector should be robust against adversarial methods. To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine authored and human authored portions might be more appropriate. Further in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open source LLM. Similarly, for AI misuse over social platforms, users are more likely to use a open-sourced model due to better flexibility and privacy. Hence, building models and benchmarks with a appropriate set of LLMs might be more applicable for practical scenarios. Many proprietary systems struggle at the task of fine-grained detection, further a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024). Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts.
3 Dataset
Our dataset consists of around 2.45M samples. We used 12 different LLMs out of which 9 are popular proprietary LLMs : GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), Amazon-Nova-Lite. We also included 3 open-source LLMs i.e Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), Mistral-large-2411 (Mistral AI, 2024) which produced outputs that are relatively difficult to distinguish from human written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets. The samples range from 30 to 25K words in length with an average length of around 600 words.
3.1 Dataset Distribution
The language distribution of the dataset and LLMs used can be seen in Figure 1. Each language-LLM pair has roughly 10000 samples. Among each set of the 10000 samples; training, development and test sets constitute 40%, 10%, 50% respectively. Additionally, among each set of 10000 samples, 10% were Completely human written, another 10% completely machine generated, and the other 80% were human-LLM co-authored i.e few portions of the text are machine generated and the rest are human written.
Figure 1: Dataset distribution per each generator and language in our dataset
3.2 Dataset Creation
GPT-4o was used through a Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/. command-r-plus and aya-23 were used through cohere’s API platform [^3] https://dashboard.cohere.com/. Rest of the models were used through open router’s [^4] https://openrouter.ai/models API. The Rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened or a rephrased version of the same text. The samples which returned the exact text or a very similar text were once again regenerated. The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text. The split locations were chosen randomly starting from the 30th word to end of text. This was done to provide the LLM with enough context to better work towards text completion.
3.3 Original Data Source and Filtering
With a goal of training on one domain and testing on every other, we chose to train on old newspapers (HC-Corpora, ) as it has sufficient number of samples i.e 17.2M for 67 languages of the same domain. We then removed samples which originated after release of gpt-3 to avoid mislabelling of samples in our dataset. Further we sampled texts which were at least 3 sentences or 50 words long. For Chinese and Japanese, we sampled texts which were at least 100 characters long.
4 Our System
We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach. We found that using additional CRF layer produced better results compared to other setups with the same model. All of the transformer models tested have produced nearly identical results over our test set. However, xlm-longformer gave better results over unseen domains and generators’ texts, and was used in the end given the longer default context length of 16384. The token level predictions by the models were then mapped into word-level predictions. We use the model’s predictions to separate text portions based on perceived authorship.
5 Evaluation and Results
We evaluate the models at 3 levels of granularity : word level, sentence level and overall. For Chinese and Japanese, we performed evaluation at a character level instead of word-level. Each domain and user might have a different preference towards metrics and evaluation, hence we report 3 metrics at each level of granularity : accuracy, recall and precision. For word level mapping of predictions, in cases where part of a word i.e a few tokens are classified differently than others, we assigned the same label to the word as its first token. While mapping word level predictions to a sentence we used majority voting, and in cases where consensus is not obtained, we assigned the same label as the first word. For evaluation over other benchmarks requiring binary classification of texts as human or machine written, we assign a human written label to the text if at least two thirds of the words get classified as human written. We also report several metrics, some of which can be seen in the below tables, rest can be found in Appendix D.
5.1 Seen Domains & Seen Generators
The results of our models over our dataset’s test set can be seen in Table 1. The samples from both the data splits are of the same domain and originate from the same set of generators.
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 97.16 90.69 97.55 96.44
Chinese* 93.13 76.28 91.40 86.58
Czech 96.23 79.63 93.84 94.98
Dutch 96.83 77.60 94.13 95.31
English 97.32 90.23 97.68 96.02
French 96.89 74.46 96.52 94.91
German 96.64 76.54 95.92 95.28
Greek 96.25 82.21 92.08 94.37
Hebrew 96.52 80.56 95.34 95.70
Hindi 97.08 92.60 97.24 96.34
Indonesian 97.20 84.92 97.19 96.64
Italian 96.44 80.69 96.84 95.38
Japanese* 92.74 83.80 92.81 86.13
Korean 97.29 84.13 94.74 95.77
Persian 96.60 88.61 96.19 94.36
Polish 96.63 88.52 92.75 95.94
Portuguese 96.46 88.51 90.29 94.89
Romanian 97.59 78.06 95.15 96.10
Russian 96.64 79.98 95.58 94.02
Spanish 96.38 71.60 96.69 94.47
Turkish 95.74 83.00 94.48 93.62
Ukrainian 95.74 74.03 96.57 93.53
Vietnamese 94.41 77.99 96.65 89.67
Average 96.26 81.94 95.11 94.19
Table 1: Word-Level Accuracy (.2f) of the models on the test dataset for each case
* Character level evaluations were done instead for Japanese and Chinese
5.2 Unseen Domains & Unseen Generators
Metrics $\rightarrow$ Accuracy Precision Recall F1
Initial Model 86.51 91.61 87.46 89.49
Final Model 86.00 87.16 92.25 89.63
Table 2: Word level Metrics over Mgtd-bench (.2f)
The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data. The benchmark consists of 11,123 samples of peer reviews and student essays (Koike et al., 2024), the generators used were various versions of llama-2 and chat-gpt (earlier version of gpt-4). the samples would hence be from completely unseen domains and generators to our models. The results of both models can be seen in Table 2.
5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
The models were tested by training on just 10k samples each from Aya-23 for English and Arabic Separately. The benchmark’s samples for Arabic were from (Alfaifi, 2013) and (Zaghouani et al., 2024). The samples for English consist of ETS and IELTS student essays sampled from non-native speakers (Chowdhury et al., 2025). Our models were used for inference directly over these texts and the strings of predicted tokens were then used to for binary classification based on how frequently the perceived authorship changed from human to LLM and vice-versa i.e the number of changes and whether the longest string consists of ones or zeroes. The metrics obtained for each language can be seen in Table 3.
Metrics $\rightarrow$ Accuracy Precision Recall F1
Arabic 95.9 96.1 94.5 95.2
English 99.1 98.7 99.3 99.0
Table 3: Overall Metrics over ETS essays (.1f)
5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
We have also tested over raid-bench (Dugan et al., 2025) which consists of texts from 11 generators and 8 domains. among them roughly 10% would be from a seen domain (news articles) while the rest are unseen by our models. The dataset’s texts were also created using various sampling strategies (greedy, random, etc.). The texts were also modified to have adversarial methods including homo-glyphs, mis-spellings, alternative spellings, article deletion etc. Among the 11 generators used, Gpt-4 is one which is similar to the generator whose outputs our model has been trained on (Gpt-4o). However, both of them have different linguistic and stylistic features, similar to how Gpt-4 is different from Gpt-3. We have tested our model’s performance once again upon being trained on our own full training data. Additionally, we have also performed an error analysis to find out what domains, models, attack strategies and decoding strategies effected the model’s performance and to what extent. This can be seen in Figure 2, Figure 3, Figure 4 and Figure 5. The texts were classified as machine generated if at least one third of the tokens within the model’s context length were classified as machine generated. The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset. Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of thse adversarial methods.
Figure 2: F1 scores VS text sampling method used
Figure 3: F1 scores VS the generator’s texts
Figure 4: F1 scores VS each domain’s texts
Figure 5: F1 scores VS adversarial method used in the input texts
6 Other Observations
The sentences inside which text authorship switches from human to LLM or vice versa were found to be relatively shorter that the original text portions which they replaced. LLMs may be likely to finish the current sentence earlier than usual to move on to the next sentence in text completion scenarios. The mean length of the original portion and the replaced portions of those sentences for each language and generator can be seen in Table 4 and Table 5 respectively. This observation was consistent across all languages and generators with a 20-30% reduction and a larger reduction in Hindi. For Chinese and Japanese too, we did observe a 20-30% reduction in character count when comparing the original and replaced portions of the sentence after the text boundary. Although there is a good variation in this feature across languages, the mean and medians observed for each language were similar for all the LLMs. This is further elaborated in Appendix A.
Language Length of Length of
Original part generated part
Arabic 17 13
Czech 11 8
Dutch 12 10
English 15 11
French 14 11
German 12 9
Greek 15 12
Hebrew 11 9
Hindi 26 12
Indonesian 11 8
Italian 15 14
Korean 9 7
Persian 17 15
Polish 10 7
Portuguese 15 11
Romanian 14 11
Russian 11 9
Spanish 15 12
Turkish 10 8
Ukrainian 11 8
Vietnamese 18 14
Average 13.8 10.4
Table 4: Median length (words) of original & newly generated parts of the sentences : Language wise
Generator Length of Length of
original part generated part
Amazon-Nova-Pro 14 10
Amazon-Nova-Lite 12 10
Aya-23-35B 11 10
Claude-3.5-Haiku 18 10
Claude-3.5-Sonnet 16 10
Command-R-Plus 16 10
GPT-4o 12 10
GPT-o1 11 9
Gemini-1.5-Pro 15 10
Gemini-1.5-Flash 9 10
Mistral-Large-2411 11 10
Perplexity-Sonar-large 15 11
Average 13.3 10
Table 5: Median length (words) of original & newly generated parts of the sentences : Generator wise
Figure 6: Accuracy VS length of input texts (sentence count)
Generator Accuracy
Amazon-Nova-Pro 94.90
Amazon-Nova-Lite 95.26
Aya-23-35B 91.75
Claude-3.5-Haiku 96.07
Claude-3.5-Sonnet 95.97
Command-R-Plus 93.92
GPT-4o 91.78
GPT-o1 96.61
Gemini-1.5-Flash 92.34
Gemini-1.5-pro 93.38
Mistral-Large-2411 93.47
Perplexity-Sonar-large 94.91
Average 94.31
Table 6: Word level accuracy (.2f) of our models over our dataset (English)
* excluding Chinese and Japanese
7 Conclusion
Despite not being trained on the domains or generators, the models built through our approach performed well over several benchmarks as seen in subsection 5.3 and subsection 5.4 over inputs which were from non-native speakers and consist of adversarial methods. Further, one case where many proprietary systems struggle is when the inputs were too short, which our models were able to overcome as seen in Figure 6, which demonstrates our models’ accuracy over our test set compared to input text’s sentence count. Table 6 displays our model’s performance over English subset of our dataset for each generator. A similar trend from subsection 5.4 was observed with models which are likely less instruction-tuned / not instruction-tuned tend to produce texts which are harder to distinguish than their alternatives.
7.1 Scalability and scope for extension
The original dataset used to train our current models as mentioned in section 3 consists of samples over 60 languages which would cover 70% of the world population’s primary language, and all of the languages are supported by existing multilingual transformer models making the process of scaling the work to more languages easier. Despite not being trained on the generators or domains’ texts, our models were able to perform well on several benchmarks. Even reaching a F1 score of 0.79 against adversarial inputs while they were neither trained over them nor pre-processed. Similarly, creation and usage of such large datasets of other domains along with ours might result in robust and better models. We couldn’t explore the relation between instruction tuning sample size of LLMs and detectability of their texts due to the proprietary nature of most of the generators we used, but a similar study using open-data models could uncover more insights.
7.2 Scope for Improvement
As seen in Figure 5, almost none of the adversarial methods affected the models built through our approach other than paraphrasing and homo-glyphs. However homo-glyphs can be pre-processed by mapping them to the actual character they were imitating in the text. This would require a large collection of homo-glyph to character mapping set to use for pre-processing. Further, paraphrased samples of various number of iterations being included in the training dataset might lead to further improvements. It is also worth exploring how detectable are texts in cases where multiple generators contribute a portion each in a human authored text. Other missing adversarial methods that are likely to be used in practical scenarios include usage of proprietary systems that ’humanize’ a given text in an attempt to evade detection.
7.3 Ideal Usage
The models were built primarily for a human-in-the-loop use cases where the model would try to flag most of the likely machine-generated portions while the flagged content can be validated either through an ensemble of models or a human and hence a tilt towards higher recall can be observed in the metrics as seen in Table 12.
Limitations
Just like any other detector or classifier, no detector can guarantee a 100% accuracy and hence the models are not meant to be used directly for decision making but are meant to be used in a human-in-the-loop scenarios. Furthermore, the experiments carried out did not include cases of multiple LLMs co-authoring a portion each of the same text.
Appendix A Pre- and Post- Boundary Comparisons
The mean and median word counts of the text portions in a sentence after the text authorship shifts from human to LLM can be seen in Table 10 and Table 11 in comparison to the texts they replace.
Appendix B Dataset Creation
The max_new_tokens value specified to the generator during creation of partial cases was randomized between 80% to 200% of the length of the portion that is being replaced. The prompts used for creation of the partial samples and rewritten samples can be seen in Table 7 and Table 8 respectively.
continue this text in Language directly :
complete this text in Language, respond directly :
Table 7: Prompts used in dataset creation : Partial cases
Rewrite this in Language a different way :
Generate an alternative version of this in Language :
Generate a later update to this in Language :
Generate a previous version of this in Language ;
Table 8: Prompts used in dataset creation : Rewritten cases
Hyperparameter Value
Seed (Training) 1024
Seed (Shuffling) 1024
Number of Epochs 5
Per Device Batch Size (Train) 12
Per Device Batch Size (Eval) 30
Context Length 16384
Learning Rate 5e-5
Weight Decay 0
Dropout (CRF Layer) 0.075
Table 9: Training Hyper-parameters used
Language $\downarrow$ Mean length of Mean length of Median length of Median length of
old text portion new text portion Old text portion New text portion
Arabic 18.73 16.25 17 13
Czech 12.02 9.52 11 8
Dutch 13.73 13.39 12 10
English 16.04 14.59 15 11
French 15.50 13.16 14 11
German 13.03 10.89 12 9
Greek 16.74 14.87 15 12
Hebrew 12.64 10.65 11 9
Hindi 40.56 15.42 26 12
Indonesian 12.44 9.56 11 8
Italian 17.54 16.39 15 14
Korean 9.85 8.08 9 7
Persian 18.83 19.88 17 15
Polish 11.42 8.84 10 7
Portuguese 16.52 13.29 15 11
Romanian 16.30 13.50 14 11
Russian 12.27 10.63 11 9
Spanish 17.18 14.81 15 12
Turkish 11.81 9.74 10 8
Ukrainian 12.04 10.39 11 8
Vietnamese 20.06 18.01 18 14
Average 16.19 12.95 13.76 10.43
Table 10: Comparison of replaced and generated text portion lengths (word count) : Language wise
[…]
old text portion new text portion Old text portion New text portion
Amazon-Nova-Pro 16.02 13.27 12 10
Amazon-Nova-Lite 18.12 12.87 14 10
Aya-23-35B 13.70 12.87 11 10
Claude-3.5-Haiku 20.19 13.13 18 10
Claude-3.5-Sonnet 17.32 12.98 16 10
Command-R-Plus 16.92 13.28 16 10
GPT-4o 13.49 12.84 12 10
GPT-o1 14.83 12.36 11 9
Gemini-1.5-Flash 19.68 13.44 15 10
Gemini-1.5-pro 12.50 13.48 9 10
Mistral-Large-2411 12.85 12.85 11 10
Perplexity-Sonar-large 17.13 13.45 15 11
Average 16.06 13.07 13.33 10
Table 11: Comparison of replaced and generated text portion lengths (word count) : Generator wise
Appendix C Reproducibility
We used multilingual longformer  [^5] https://huggingface.co/hyperonym/xlm-roberta-longformer-base-16384 with an additional CRF layer. The hyper-parameters used for training the models can be seen in Table 9. We built a separate model for each language, the training was done over A100 SXM over 10h each.
Appendix D Other Metrics
The metrics over each type of text for each language and LLM separately can be seen in Table 13, Table 14, Table 15, Table 16, Table 17, Table 18, Table 19, Table 20, Table 21, Table 22, Table 23, Table 24.
Appendix E License
The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0666https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.
Language $\downarrow$ Accuracy Precision Recall F1-score
Arabic 96.44 92.50 97.17 94.78
Chinese* 86.58 87.03 86.46 86.75
Czech 94.98 94.57 97.96 96.23
Dutch 95.31 93.34 97.97 95.60
English 96.02 92.34 98.44 95.29
French 94.91 93.64 98.42 95.97
German 95.28 94.87 98.38 96.59
Greek 94.37 93.69 96.51 95.08
Hebrew 95.70 95.32 97.94 96.61
Hindi 96.34 89.72 96.66 93.06
Indonesian 96.64 95.61 98.29 96.93
Italian 95.38 95.04 97.58 96.29
Japanese* 86.13 85.64 94.17 89.70
Korean 95.77 95.29 97.69 96.48
Persian 94.36 84.45 96.88 90.24
Polish 95.94 96.76 97.19 96.97
Portuguese 94.89 91.92 96.07 93.95
Romanian 96.10 95.81 98.53 97.15
Russian 94.02 86.67 97.29 91.67
Spanish 94.47 90.02 98.14 93.90
Turkish 93.62 88.56 97.17 92.66
Ukrainian 93.53 86.58 97.93 91.90
Vietnamese 89.67 77.23 97.44 86.17
Average 94.19 91.16 96.97 93.91
Table 12: Word-level Metrics of our models over each language : our test set
* Character level evaluations were done instead for Japanese and Chinese
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 97.56 88.10 98.09 97.17
Chinese* 93.70 75.35 91.60 87.00
Czech 96.63 80.10 94.36 95.20
Dutch 95.21 78.00 92.10 95.23
English 97.77 89.61 98.87 96.60
French 97.34 72.85 97.14 95.28
German 96.92 75.73 95.29 95.58
Greek 95.65 81.80 82.85 92.96
Hebrew 97.35 70.89 96.26 95.73
Hindi 96.65 92.82 96.67 96.59
Indonesian 97.27 85.73 95.76 96.60
Italian 96.88 80.88 94.70 95.65
Japanese* 97.48 88.57 93.94 96.85
Korean 97.68 84.15 93.73 95.39
Persian 96.91 89.45 93.22 94.05
Polish 96.96 87.52 92.32 95.98
Portuguese 95.28 94.15 96.32 95.28
Romanian 96.53 76.55 96.64 96.23
Russian 96.46 79.10 94.45 94.03
Spanish 96.92 71.75 96.97 94.97
Turkish 95.55 82.68 98.17 92.65
Ukrainian 95.39 73.81 95.48 93.90
Vietnamese 94.46 76.17 97.14 88.74
Table 13: Case wise accuracies over all languages for each generator : amazon-nova-pro
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 96.56 90.93 95.62 95.88
Chinese* 93.20 76.99 93.57 87.77
Czech 97.81 79.40 94.58 96.43
Dutch 97.07 78.10 92.85 95.19
English 98.11 89.25 98.73 96.80
French 97.59 76.28 97.79 96.07
German 98.01 76.34 95.52 96.76
Greek 96.00 79.78 88.01 93.66
Hebrew 98.05 83.84 94.35 96.78
Hindi 96.49 91.59 95.30 95.38
Indonesian 97.99 85.03 97.18 97.06
Italian 96.95 80.81 95.45 95.54
Japanese* 98.07 76.50 93.02 92.78
Korean 98.20 82.49 95.42 95.68
Persian 97.40 88.48 94.92 95.31
Polish 97.55 89.32 93.83 96.63
Portuguese 92.67 87.92 95.09 94.35
Romanian 97.97 76.44 93.76 96.20
Russian 97.22 81.47 96.30 95.10
Spanish 97.49 71.55 97.15 94.98
Turkish 96.63 83.99 90.84 93.87
Ukrainian 96.74 74.80 99.91 94.24
Vietnamese 95.28 78.87 97.03 89.76
Table 14: Case wise accuracies over all languages for each generator : amazon-nova-lite
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 95.14 92.22 97.76 96.05
Chinese* 85.65 75.66 89.10 82.40
Czech 89.75 79.02 87.94 89.24
Dutch 92.45 79.97 92.99 93.12
English 93.01 90.49 96.96 93.52
French 93.28 73.12 95.14 92.72
German 89.89 77.28 92.53 90.06
Greek 92.04 80.69 91.83 91.75
Hebrew 96.71 82.75 91.54 95.32
Hindi 94.18 93.62 92.96 94.88
Indonesian 90.91 83.55 95.28 92.85
Italian 89.21 75.43 88.87 88.87
Japanese* 75.56 78.10 91.21 75.64
Korean 95.04 85.46 92.92 94.14
Persian 93.81 87.28 95.29 92.98
Polish 90.40 86.41 89.15 90.81
Portuguese 92.69 91.17 90.96 92.69
Romanian 93.65 78.15 95.16 93.17
Russian 93.00 79.77 92.12 92.20
Spanish 91.30 72.87 93.17 91.88
Turkish 90.19 82.59 98.19 90.77
Ukrainian 87.77 73.69 97.57 90.69
Vietnamese 87.70 76.08 96.83 88.54
Table 15: Case wise accuracies over all languages for each generator : Aya-23
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 98.82 91.63 95.75 97.18
Chinese* 86.51 75.93 86.18 87.75
Czech 99.80 80.78 91.56 97.55
Dutch 99.49 77.57 86.16 96.41
English 99.37 90.98 98.32 97.76
French 99.63 74.86 90.15 96.30
German 99.79 77.23 94.62 97.37
Greek 99.90 87.33 82.84 97.46
Hebrew 98.94 83.48 82.61 96.27
Hindi 98.72 92.35 95.48 97.23
Indonesian 99.55 88.19 94.11 98.05
Italian 99.97 81.43 93.49 97.48
Japanese* 98.33 87.97 91.08 97.02
Korean 99.40 84.22 94.45 96.93
Persian 97.99 89.54 90.00 94.54
Polish 99.60 88.75 85.69 97.62
Portuguese 99.17 90.82 82.19 96.52
Romanian 99.93 78.70 92.11 97.11
Russian 99.26 80.44 92.17 95.36
Spanish 99.31 71.65 93.24 95.91
Turkish 98.60 81.65 92.86 94.62
Ukrainian 99.27 73.52 91.46 93.99
Vietnamese 98.42 77.05 92.41 91.85
Table 16: Case wise accuracies over all languages for each generator : Claude-3.5-Haiku
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 98.63 91.82 100.00 96.58
Chinese* 92.64 78.36 95.14 88.43
Czech 99.30 77.22 99.88 97.62
Dutch 99.30 77.53 99.52 97.30
English 99.35 90.02 99.76 98.03
French 99.53 73.66 99.97 97.06
German 99.52 76.06 99.69 97.33
Greek 99.00 80.60 99.62 95.83
Hebrew 97.68 82.69 99.88 96.46
Hindi 99.12 92.51 99.88 97.63
Indonesian 99.66 84.55 100.00 98.43
Italian 99.69 81.13 99.99 98.13
Japanese* 98.59 87.36 99.64 98.04
Korean 98.77 83.49 99.87 97.15
Persian 98.35 87.92 99.97 96.01
Polish 99.00 90.30 99.26 98.20
Portuguese 98.74 89.65 83.74 96.38
Romanian 99.18 80.36 99.71 97.46
Russian 99.33 80.55 99.93 94.55
Spanish 99.06 71.68 99.92 96.65
Turkish 98.45 83.13 99.96 95.32
Ukrainian 99.06 74.14 99.87 95.62
Vietnamese 98.10 77.40 99.92 88.87
Table 17: Case wise accuracies over all languages for each generator : Claude-3.5-Sonnet
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 87.12 85.34 86 82
Chinese* 88.90 86.45 89 84
English 92.45 90.12 91 88
French 89.78 87.21 90 85
German 90.23 88.05 89 86
Italian 89.12 87.00 89 86
Japanese* 87.77 85.88 88 83
Korean 88.56 86.34 87 85
Portuguese 90.12 88.34 89 87
Spanish 90.45 88.12 89 87
Table 18: Case wise accuracies over all languages for each generator : Command-R-Plus
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 95.74 91.26 96.31 95.01
Chinese* 92.63 77.87 92.51 86.36
Czech 93.30 80.96 91.67 91.87
Dutch 94.14 74.85 90.84 92.58
English 94.94 90.02 92.84 94.01
French 92.87 75.18 93.90 90.89
German 93.67 75.82 93.99 91.97
Greek 94.22 81.18 95.67 92.11
Hebrew 92.60 82.51 95.10 91.85
Hindi 96.56 92.44 96.95 96.16
Indonesian 95.08 84.88 95.88 94.70
Italian 93.35 79.95 92.72 92.72
Japanese* 93.98 88.44 94.19 93.84
Korean 94.44 84.84 93.09 93.61
Persian 94.83 88.32 94.68 93.34
Polish 94.53 89.51 89.36 93.73
Portuguese 95.50 88.58 85.07 93.93
Romanian 94.59 77.44 92.73 93.26
Russian 92.90 80.17 97.34 92.61
Spanish 93.54 69.64 91.87 92.13
Turkish 92.83 83.80 88.09 91.37
Ukrainian 91.69 74.93 96.81 90.33
Vietnamese 92.10 77.39 93.32 88.25
Table 19: Case wise accuracies over all languages for each generator : GPT-4o
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 99.10 90.46 97.08 97.92
Chinese* 95.08 76.01 86.18 87.30
Czech 98.84 80.76 86.30 97.05
Dutch 98.80 77.47 89.03 97.12
English 99.07 88.92 94.25 96.91
French 98.77 76.17 91.74 96.53
German 98.92 76.66 89.69 97.12
Greek 98.87 81.60 85.68 97.05
Hebrew 98.97 83.94 97.33 98.10
Hindi 99.10 92.12 97.42 97.33
Indonesian 98.96 84.98 99.00 98.45
Italian 97.04 80.93 99.10 96.16
Japanese* 90.78 73.63 85.24 78.08
Korean 99.16 83.18 83.13 97.09
Persian 98.72 87.01 94.43 95.50
Polish 99.04 90.29 86.92 97.86
Portuguese 98.65 88.47 83.50 95.18
Romanian 98.77 76.50 98.37 97.57
Russian 98.98 78.13 87.02 95.23
Spanish 98.80 71.70 92.79 96.12
Turkish 98.94 82.37 87.33 96.55
Ukrainian 99.05 75.07 93.34 96.45
Vietnamese 98.29 78.87 91.75 92.24
Table 20: Case wise accuracies over all languages for each generator : GPT-o1
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 93.90 88.40 98.59 94.86
Chinese* 89.29 75.98 94.92 85.09
Czech 91.43 78.82 97.15 92.04
Dutch 95.08 78.10 91.91 94.86
English 96.57 91.03 97.34 94.64
French 95.92 76.69 98.72 94.04
German 96.22 78.67 98.80 95.15
Greek 92.04 84.32 98.56 93.52
Hebrew 91.90 69.05 98.50 92.51
Hindi 95.55 93.52 98.72 95.66
Indonesian 96.84 83.45 98.55 95.92
Italian 94.70 76.47 97.03 94.89
Japanese* 84.59 87.53 96.26 87.86
Korean 94.54 85.07 99.41 94.66
Persian 95.68 89.47 96.54 94.66
Polish 93.51 88.07 97.34 94.60
Portuguese 95.22 88.75 94.90 94.28
Romanian 97.31 75.73 97.53 96.30
Russian 94.96 78.53 99.40 92.82
Spanish 95.30 74.56 99.47 94.09
Turkish 94.82 82.44 96.67 92.51
Ukrainian 89.37 73.96 98.39 91.68
Vietnamese 91.06 80.25 99.45 87.71
Table 21: Case wise accuracies over all languages for each generator : Gemini-1.5-Pro
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 98.55 89.55 98.00 95.88
Chinese* 89.94 75.81 94.03 87.75
Czech 98.45 80.42 99.78 96.67
Dutch 98.22 78.00 99.39 95.86
English 97.92 90.02 99.39 95.26
French 98.10 73.66 99.94 95.59
German 98.71 74.75 99.58 96.74
Greek 98.96 85.16 98.90 98.09
Hebrew 97.64 82.90 99.71 96.82
Hindi 98.17 93.06 99.62 97.32
Indonesian 99.11 85.64 99.72 97.65
Italian 98.25 83.62 99.89 98.28
Japanese* 95.71 88.47 97.35 95.88
Korean 98.36 84.40 99.36 97.10
Persian 96.67 88.45 95.52 93.87
Polish 99.01 87.67 99.27 97.84
Portuguese 96.72 88.12 90.12 95.52
Romanian 99.84 77.30 99.75 98.49
Russian 97.62 81.30 99.07 94.29
Spanish 97.50 68.98 99.91 94.37
Turkish 97.69 84.99 99.36 95.47
Ukrainian 97.46 75.89 99.67 93.70
Vietnamese 96.29 79.35 99.81 91.12
Table 22: Case wise accuracies over all languages for each generator : Gemini-1.5-Flash
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
Arabic 96.53 92.00 99.40 95.68
Chinese* 95.05 76.30 97.63 88.07
Czech 96.99 78.78 94.73 94.98
Dutch 94.46 78.10 91.01 94.46
English 96.74 90.32 99.80 95.91
French 97.06 74.65 97.22 94.35
German 96.69 76.14 98.43 94.77
Greek 95.78 79.75 96.39 92.69
Hebrew 95.40 83.36 97.30 93.86
Hindi 96.25 92.00 99.35 95.29
Indonesian 96.67 83.15 96.32 95.55
Italian 97.34 82.59 99.30 96.01
Japanese* 97.05 87.46 94.90 96.12
Korean 96.65 82.64 97.39 95.02
Persian 94.75 89.34 98.55 92.34
Polish 96.69 87.13 93.69 95.36
Portuguese 96.37 88.20 93.68 94.69
Romanian 97.22 77.49 97.42 95.45
Russian 96.64 80.58 97.90 93.09
Spanish 95.54 69.92 98.82 92.76
Turkish 91.99 80.61 98.47 91.99
Ukrainian 96.21 74.58 97.16 93.20
Vietnamese 92.42 78.44 98.60 88.25
Table 23: Case wise accuracies over all languages for each generator : Mistral-Large-2411
Language $\downarrow$ Partial cases Unchanged cases Rewritten cases Overall
English 97.10 91.08 99.71 96.64
French 95.53 72.58 99.49 93.94
German 94.98 77.21 99.50 94.29
Portuguese 92.66 89.06 98.17 94.03
Spanish 94.79 72.31 99.80 93.99
Table 24: Case wise accuracies over all languages for each generator : Perplexity-Sonar-Large

### v2

@@
 We feel our findings and datasets can aid in further research into mitigating the harms of AI generated texts.
+Source Dataset/Benchmark Samples Languages Generators
+(Lee et al., 2022) CoAuthor 1,445 1 1
+(Zhang et al., 2024) MixSet 3,600 1 12
+(Dugan et al., 2022) RoFT 21,646 1 5
+(Macko et al., 2024b) MultiTude 4,070 11 8
+(Artemova et al., 2025) Beemo 19,600 1 10
+Our Work placeholder 2,447,221 23 12
+Table 1: Comparison with other Human-LLM co-authored datasets & benchmarks
 2 Related Works
@@
 Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts.
+Further, the task of detection of such human-llm co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025) (Huang et al., 2025).
 3 Dataset
@@
 The samples range from 30 to 25K words in length with an average length of around 600 words.
+Table 1 provides an overview of comparison with other datasets and benchmarks for fine-grained detection of human-llm co-authored texts.
+Our dataset utilizes a better choice of generators which are more likely to be used in practical scenarios where as prior datasets are limited to smaller and limited number of models.
+Our dataset also compromises of over 100 times more samples, aiding other researchers in the field.
 3.1 Dataset Distribution
@@
 The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text.
-The split locations were chosen randomly starting from the 30th word to end of text.
+The split locations were chosen randomly from between the 30th word up to end of text.
 This was done to provide the LLM with enough context to better work towards text completion.
@@
 We use the model’s predictions to separate text portions based on perceived authorship.
+Improving the performance required balancing pre and post boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
 5 Evaluation and Results
@@
 Final Model 86.00 87.16 92.25 89.63
-Table 2: Word level Metrics over Mgtd-bench (.2f)
+Table 3: Word level Metrics over Mgtd-bench (.2f) through our models (zero-shot, unseen domains, unseen generators)
 The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data.
@@
 English 99.1 98.7 99.3 99.0
-Table 3: Overall Metrics over ETS essays (.1f)
+Arabic-Best 96.1 96.1 95.0 95.5
+English-Best 99.3 99.0 99.2 99.1
+Table 4: Overall Metrics over ETS essays (.1f) through our detectors (zero-shot, unseen generators, unseen domain) VS best submissions (fine-tuned on same generators and domain)
 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
@@
 The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset.
-Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of thse adversarial methods.
-Figure 2: F1 scores VS text sampling method used
-Figure 3: F1 scores VS the generator’s texts
-Figure 4: F1 scores VS each domain’s texts
-Figure 5: F1 scores VS adversarial method used in the input texts
+Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of the adversarial methods.
+Figure 2: F1 scores VS text sampling method used : Sampling strategy did not effect detection capability
+Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect.
+Figure 4: F1 scores VS each domain’s texts : news i.e the only training data domain was easier to detect.
+Figure 5: F1 scores VS adversarial method used in the input texts : homo-glyphs are the only real issue, and paraphrases to a small extent, while the rest can be handled through pre-processing.
+5.5 Comparisons with proprietary systems
+While many proprietary systems claim to have excellent results, they often struggle with unseen domains and generators.
+Further many systems like ZeroGPT [^5] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^6] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API.
+This would require manually evaluating the results through the UI by counting correct and incorrectly classified word counts.
+Hence, these systems were not used for comparisons for fine-grained scores.
+Several users have tried this manfully over a subset (Kadiyala, 2024) only to find a large gap in performance.
+For binary classifications as well, they have a minimum length threshold to even work, which would become incompatible over a large number of samples of the benchmark.
 6 Other Observations
@@
 Average 13.8 10.4
-Table 4: Median length (words) of original & newly generated parts of the sentences : Language wise
+Table 5: Median length (words) of original & newly generated parts of the sentences - Language wise : Models tend to finish off current sentence after authorship switch quickly before continuing with the rest of the text.
 Generator Length of Length of
@@
 Furthermore, the experiments carried out did not include cases of multiple LLMs co-authoring a portion each of the same text.
+Acknowledgements
+Several stages of this work were supported by grants from Traversaal.ai, Cohere Labs & Microsoft Azure.
 Appendix A Pre- and Post- Boundary Comparisons
@@
 The hyper-parameters used for training the models can be seen in Table 10.
-We built a separate model for each language, the training was done over A100 SXM over 10h each.
+We built a separate model for each language, the training was done over H100 SXM over 10h each.
 Appendix D Other Metrics
-The metrics over each type of text for each language and LLM separately can be seen in Table 13, Table 14, Table 15, Table 16, Table 17, Table 18, Table 19, Table 20, Table 21, Table 22, Table 23, Table 24.
+The metrics over each type of text for each language and LLM separately can be seen in Table 14 to Table 25 respectively..
 Appendix E License
-The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0666https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.
+The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0888https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.
 Language $\downarrow$ Accuracy Precision Recall F1-score

### v3

@@
 Abstract
-An ideal detection system for machine generated content is supposed to work well on any generator as many more advanced LLMs come into existence day by day.
+An ideal detection system for machine-generated content is supposed to work well on any generator as many more advanced LLMs come into existence day by day.
 Existing systems often struggle with accurately identifying AI-generated content over shorter texts.
-Further, not all texts might be entirely authored by a human or LLM, hence we focused more over partial cases i.e human-LLM co-authored texts.
-Our paper introduces a set of models built for the task of token classification which are trained on an extensive collection of human-machine co-authored texts, which performed well over texts of unseen domains, unseen generators, texts by non-native speakers and those with adversarial inputs.
-We also introduce a new dataset of over 2.4M such texts mostly co-authored by several popular proprietary LLMs over 23 languages.
-We also present findings of our models’ performance over each texts of each domain and generator.
-Additional findings include comparison of performance against each adversarial method, length of input texts and characteristics of generated texts compared to the original human authored texts.
+Further, not all texts might be entirely authored by a human or LLM; hence, we focused more on partial cases, i.e., human-LLM co-authored texts.
+Our paper introduces a set of models built for the task of token classification that are trained on an extensive collection of human-machine co-authored texts, which performed well over texts of unseen domains, unseen generators, texts by non-native speakers, and those with adversarial inputs.
+We also introduce a new dataset of over 2.4M such texts, mostly co-authored by several popular proprietary LLMs in over 23 languages.
+We also present findings of our models’ performance over texts of each domain and generator.
+Additional findings include a comparison of performance against each adversarial method, the length of input texts, and the characteristics of generated texts compared to the original human-authored texts.
 1 Introduction
@@
 However, as LLMs become more sophisticated, existing detection systems must be re-evaluated and enhanced to maintain their effectiveness.
-Further, Each domain comes with its version of the issue of detecting machine generated texts.
+Further, each domain comes with its version of the issue of detecting machine-generated texts.
 For instance, proprietary LLMs with internet access and better knowledge cutoffs are more likely to be used in domains like academia.
-Similarly, bad actors might use an open source generators for the task of creating misinformation and deception through machine generated online content as such models can be hosted locally to not leave a trail and are more flexible in terms of not denying user requests.
-Hence, tailoring models and approaches for each specific domain/scenario might be better applicable for practical scenarios.
-We chose a token-classification approach to train a model for the task of distinguish writing styles within a text if more than one were found.
-This approach helped us achieve better performance over texts of unseen features (i.e domain, generator, adversarial inputs, non-native speakers’ texts) as our models were trained to distinguish different styles within a text rather than classifying an input text as one of the two classes it was trained on.
-Further, we explored the findings and results upon testing our models over other benchmarks which consist of texts from unseen domain and generators.
-We also tested our models over benchmarks which consist of texts with various adversarial inputs and those written by non-native speakers.
-We feel our findings and datasets can aid in further research into mitigating the harms of AI generated texts.
+Similarly, bad actors might use an open-source generator for the task of creating misinformation and deception through machine-generated online content, as such models can be hosted locally to not leave a trail and are more flexible in terms of not denying user requests.
+Hence, tailoring models and approaches for each specific domain/scenario might be more applicable for practical scenarios.
+We chose a token-classification approach to train a model for the task of distinguishing writing styles within a text if more than one was found.
+This approach helped us achieve better performance over texts of unseen features (i.e., domain, generator, adversarial inputs, and non-native speakers’ texts) as our models were trained to distinguish different styles within a text rather than classifying an input text as one of the two classes it was trained on.
+Further, we explored the findings and results upon testing our models over other benchmarks, which consist of texts from unseen domains and generators.
+We also tested our models over benchmarks, which consist of texts with various adversarial inputs and those written by non-native speakers.
 Source Dataset/Benchmark Samples Languages Generators
@@
 (Macko et al., 2024b) MultiTude 4,070 11 8
+(Wang et al., 2024c) M4GTD 31,893 1 4
 (Artemova et al., 2025) Beemo 19,600 1 10
@@
 Existing methods often struggle with accurately identifying AI-generated content over shorter texts.
-Moreover, binary classification approaches, which categorize texts as either human or AI-generated (Wang et al., 2024a), (Wang et al., 2024b), (Bhattacharjee et al., 2023), (Zellers et al., 2019b), (Macko et al., 2023), (Ghosal et al., 2023), (Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs.
-In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a), (Ignat et al., 2024), where content typically consists of one or two sentences.
-Additionally, some detection works rely on detecting watermarks from AI-generated texts, (Chang et al., 2024), (Dathathri et al., 2024), (Sadasivan et al., 2024), (Zhao et al., 2023) but not all generators utilize watermarking limiting the applicability of such approaches.
-Few other approaches utilize statistical methods (Mitchell et al., 2023), (Kumarage et al., 2023), (Gehrmann et al., 2019b), (Hans et al., 2024), (Bao et al., 2023), but they can be prone to mis-classification against adversarial methods like rephrasing and humanizing.
-(Abassy et al., 2024) introduced a 4-way classification as entirely human authored, entirely llm authored, human-edited and llm-authored or llm-edited human-authored.
+Moreover, binary classification approaches, which categorize texts as either human or AI-generated, (Wang et al., 2024a, b; Bhattacharjee et al., 2023; Zellers et al., 2019b; Macko et al., 2023; Ghosal et al., 2023; Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs.
+In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a; Ignat et al., 2024), where content typically consists of one or two sentences.
+Additionally, some detection works rely on detecting watermarks from AI-generated texts.
+(Chang et al., 2024; Dathathri et al., 2024; Sadasivan et al., 2024; Zhao et al., 2023) but not all generators utilize watermarking, limiting the applicability of such approaches.
+A few other approaches utilize statistical methods (Mitchell et al., 2023; Kumarage et al., 2023; Gehrmann et al., 2019b; Hans et al., 2024; Bao et al., 2023), but they can be prone to misclassification against adversarial methods like rephrasing and humanizing.
+(Abassy et al., 2024) introduced a 4-way classification as entirely human-authored, entirely LLM-authored, human-edited and LLM-authored, or LLM-edited human-authored.
 An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts.
-Further an ideal detector should be robust against adversarial methods.
-To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine authored and human authored portions might be more appropriate.
-Further in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open source LLM.
-Similarly, for AI misuse over social platforms, users are more likely to use a open-sourced model due to better flexibility and privacy.
-Hence, building models and benchmarks with a appropriate set of LLMs might be more applicable for practical scenarios.
-Many proprietary systems struggle at the task of fine-grained detection, further a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024).
-Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts.
-Further, the task of detection of such human-llm co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025) (Huang et al., 2025).
+Further, an ideal detector should be robust against adversarial methods.
+To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine-authored and human-authored portions might be more appropriate.
+Further, in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open-source LLM.
+Similarly, for AI misuse over social platforms, users are more likely to use an open-source model due to better flexibility and privacy.
+Hence, building models and benchmarks with an appropriate set of LLMs might be more applicable for practical scenarios.
+Many proprietary systems struggle at the task of fine-grained detection; further, a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024).
+Previous works in the similar direction include (Lee et al., 2022; Zhang et al., 2024; Dugan et al., 2023; Macko et al., 2024b; Liang et al., 2024; Chen et al., 2023), which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts.
+Further, the task of detection of such human-LLM co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025; Huang et al., 2025).
 3 Dataset
 Our dataset consists of around 2.45 M samples.
-We used 12 different LLMs out of which 9 are popular proprietary LLMs : GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), Amazon-Nova-Lite.
-We also included 3 open-source LLMs i.e Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), Mistral-large-2411 (Mistral AI, 2024) which produced outputs that are relatively difficult to distinguish from human written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets.
-The samples range from 30 to 25K words in length with an average length of around 600 words.
+We used 12 different LLMs, out of which 9 are popular proprietary LLMs: GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), and Amazon-Nova-Lite.
+We also included 3 open-source LLMs, i.e., Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), and Mistral-large-2411 (Mistral AI, 2024), which produced outputs that are relatively difficult to distinguish from human-written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets.
+The samples range from 30 to 25,000 words in length with an average length of around 600 words.
 Table 1 provides an overview of comparison with other datasets and benchmarks for fine-grained detection of human-LLM co-authored texts.
-Our dataset utilizes a better choice of generators which are more likely to be used in practical scenarios where as prior datasets are limited to smaller and limited number of models.
-Our dataset also compromises of over 100 times more samples, aiding other researchers in the field.
+Our dataset utilizes a better choice of generators, which are more likely to be used in practical scenarios, whereas prior datasets are limited to smaller and a limited number of models.
+Our dataset also comprises the largest collection of human-LLM co-authored texts, aiding other researchers in the field.
 3.1 Dataset Distribution
@@
 Each language-LLM pair has roughly 10000 samples.
-Among each set of the 10000 samples; training, development and test sets constitute 40%, 10%, 50% respectively.
-Additionally, among each set of 10000 samples, 10% were Completely human written, another 10% completely machine generated, and the other 80% were human-LLM co-authored i.e few portions of the text are machine generated and the rest are human written.
+Among each set of the 10000 samples, training, development, and test sets constitute 40%, 10%, and 50%, respectively.
+Additionally, among each set of 10000 samples, 10% were completely human written, another 10% completely machine generated, and the other 80% were human-LLM co-authored, i.e., a few portions of the text are machine generated and the rest are human written.
 Figure 1: Dataset distribution per each generator and language in our dataset
 3.2 Dataset Creation
-GPT-4o was used through a Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/.
+GPT-4o was used through an Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/.
 Command-r-plus and Aya-23 were used through Cohere’s API platform [^3] https://dashboard.cohere.com/.
-Rest of the models were used through open router’s [^4] https://openrouter.ai/models API.
-The Rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened or a rephrased version of the same text.
-The samples which returned the exact text or a very similar text were once again regenerated.
-The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text.
-The split locations were chosen randomly from between the 30th word up to end of text.
+The rest of the models were used through OpenRouter’s [^4] https://openrouter.ai/models API.
+The rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened, or a rephrased version of the same text.
+The samples that returned the exact text or a very similar text were once again regenerated.
+The partially machine-generated texts were created by splitting the text at random locations, and the generator was asked to finish the text.
+The split locations were chosen randomly from between the 30th word up to the end of the text.
 This was done to provide the LLM with enough context to better work towards text completion.
 3.3 Original Data Source and Filtering
-With a goal of training on one domain and testing on every other, we chose to train on old newspapers (HC-Corpora, ) as it has sufficient number of samples i.e 17.2M for 67 languages of the same domain.
-We then removed samples which originated after release of gpt-3 to avoid mislabelling of samples in our dataset.
-Further we sampled texts which were at least 3 sentences or 50 words long.
-For Chinese and Japanese, we sampled texts which were at least 100 characters long.
+With a goal of training on one domain and testing on every other, we chose to train on old newspapers (HC-Corpora, ) as it has a sufficient number of samples, i.e., 17.2 M for 67 languages of the same domain.
+We then removed samples that originated after the release of GPT-3 to avoid mislabeling of samples in our dataset.
+Further, we sampled texts that were at least 3 sentences or 50 words long.
+For Chinese and Japanese, we sampled texts that were at least 100 characters long.
 4 Our System
 We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach.
-We found that using additional CRF layer produced better results compared to other setups with the same model.
+We found that using an additional CRF layer produced better results compared to other setups with the same model.
 All of the transformer models tested have produced nearly identical results over our test set.
-However, XLM-longformer gave better results over unseen domains and generators’ texts, and was used in the end given the longer default context length of 16384.
-The token level predictions by the models were then mapped into word-level predictions.
+However, XLM-Longformer gave better results over unseen domains and generators’ texts and was used in the end given the longer default context length of 16384.
+The token-level predictions by the models were then mapped into word-level predictions.
 We use the model’s predictions to separate text portions based on perceived authorship.
-Improving the performance required balancing pre and post boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
+Improving the performance required balancing pre- and post-boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
 5 Evaluation and Results
-We evaluate the models at 3 levels of granularity : word level, sentence level and overall.
-For Chinese and Japanese, we performed evaluation at a character level instead of word-level.
-Each domain and user might have a different preference towards metrics and evaluation, hence we report 3 metrics at each level of granularity : accuracy, recall and precision.
-For word level mapping of predictions, in cases where part of a word i.e a few tokens are classified differently than others, we assigned the same label to the word as its first token.
-While mapping word level predictions to a sentence we used majority voting, and in cases where consensus is not obtained, we assigned the same label as the first word.
-For evaluation over other benchmarks requiring binary classification of texts as human or machine written, we assign a human written label to the text if at least two thirds of the words get classified as human written.
-We also report several metrics, some of which can be seen in the below tables, rest can be found in Appendix D.
+We evaluate the models at 3 levels of granularity: word level, sentence level, and overall.
+For Chinese and Japanese, we performed evaluation at a character level instead of a word level.
+[^5] Unlike English, Chinese and Japanese are written without spaces between words, so segmenting text into words requires an external lexicon or trained segmenter and can introduce alignment errors.
+Using characters as the atomic units avoids these segmentation ambiguities and ensures every unit is evaluated consistently.
+Each domain and user might have a different preference towards metrics and evaluation; hence, we report 3 metrics at each level of granularity: accuracy, recall, and precision.
+For word-level mapping of predictions, in cases where part of a word, i.e., a few tokens, are classified differently than others, we assign the same label to the word as its first token.
+While mapping word-level predictions to a sentence, we used majority voting, and in cases where consensus was not obtained, we assigned the same label as the first word.
+For evaluation over other benchmarks requiring binary classification of texts as human or machine written, we assign a human-written label to the text if at least one percent of the words get classified as human-written.
+We also report several metrics, some of which can be seen in the below tables; the rest can be found in Appendix D.
 5.1 Seen Domains & Seen Generators
@@
 Table 3: Word level Metrics over Mgtd-bench (.2f) through our models (zero-shot, unseen domains, unseen generators)
-The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data.
-The benchmark consists of 11,123 samples of peer reviews and student essays (Koike et al., 2024), the generators used were various versions of llama-2 and chat-gpt (earlier version of gpt-4).
-the samples would hence be from completely unseen domains and generators to our models.
-The results of both models can be seen in Table 3.
-5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
-The models were tested by training on just 10k samples each from Aya-23 for English and Arabic Separately.
-The benchmark’s samples for Arabic were from (Alfaifi, 2013) and (Zaghouani et al., 2024).
-The samples for English consist of ETS and IELTS student essays sampled from non-native speakers (Chowdhury et al., 2025).
-Our models were used for inference directly over these texts and the strings of predicted tokens were then used to for binary classification based on how frequently the perceived authorship changed from human to LLM and vice-versa i.e the number of changes and whether the longest string consists of ones or zeroes.
-The metrics obtained for each language can be seen in Table 4.
 Metrics $\rightarrow$ Accuracy Precision Recall F1
@@
 Table 4: Overall Metrics over ETS essays (.1f) through our detectors (zero-shot, unseen generators, unseen domain) VS best submissions (fine-tuned on same generators and domain)
+The models were tested twice over (Wang et al., 2024a): once by training on just 10000 samples of a single generator (Aya-23) and again later by training over our complete training data.
+The benchmark consists of 11,123 samples of peer reviews and student essays (Koike et al., 2024); the generators used were various versions of Llama-2 and ChatGPT (an earlier version of GPT-4).
+The samples would hence be from completely unseen domains and generators to our models.
+The results of both models can be seen in Table 3.
+5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
+The models were tested by training on just 10,000 samples each from Aya-23 for English and Arabic separately.
+The benchmark’s samples for Arabic were from (Alfaifi, 2013) and (Zaghouani et al., 2024).
+The samples for English consist of ETS and IELTS student essays sampled from non-native speakers (Chowdhury et al., 2025).
+Our models were used for inference directly over these texts, and the strings of predicted tokens were then used for binary classification based on how frequently the perceived authorship changed from human to LLM and vice versa, i.e., the number of changes and whether the longest string consists of ones or zeroes.
+The metrics obtained for each language can be seen in Table 4.
+System Source M E R Avg
+GLTR (Gehrmann et al., 2019b) 0.706 0.943 0.773 0.807
+Binoculars (Hans et al., 2024) 0.822 0.995 0.613 0.778
+RADAR (Gu et al., 2025) 0.768 0.939 0.716 0.794
+Current work Our Baselines 0.713 0.914 0.811 0.812
+Table 5: Comparisons with other systems on the same benchmarks: MGTD-Bench (M), Academic-Essays (E), and Raid-Bench (R) through F1 Scores, and the average scores from the 3 benchmarks (AVG).
+The results cover a randomly sampled subset of each benchmark spanning over 13k samples, including those with adversarial methods (R) and non-native speakers (E).
+Randomly sampled subsets were used with equal distribution of all generators, domains, and adversarial methods distribution in the subset for each benchmark.
 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
-We have also tested over raid-bench (Dugan et al., 2025) which consists of texts from 11 generators and 8 domains.
-among them roughly 10% would be from a seen domain (news articles) while the rest are unseen by our models.
+We have also tested over raid-bench (Dugan et al., 2025), which consists of texts from 11 generators and 8 domains.
+Among them, roughly 10% would be from a seen domain (news articles), while the rest are unseen by our models.
 The dataset’s texts were also created using various sampling strategies (greedy, random, etc.).
-The texts were also modified to have adversarial methods including homo-glyphs, mis-spellings, alternative spellings, article deletion etc.
-Among the 11 generators used, Gpt-4 is one which is similar to the generator whose outputs our model has been trained on (Gpt-4o).
+The texts were also modified to have adversarial methods, including homoglyphs, misspellings, alternative spellings, article deletion, etc.
+Among the 11 generators used, GPT-4 is one that is similar to the generator whose outputs our model has been trained on (GPT-4o).
 However, both of them have different linguistic and stylistic features, similar to how GPT-4 is different from GPT-3.
 We have tested our model’s performance once again upon being trained on our own full training data.
-Additionally, we have also performed an error analysis to find out what domains, models, attack strategies and decoding strategies effected the model’s performance and to what extent.
-This can be seen in Figure 2, Figure 3, Figure 4 and Figure 5.
-The texts were classified as machine generated if at least one third of the tokens within the model’s context length were classified as machine generated.
-The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset.
-Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of the adversarial methods.
+Additionally, we have also performed an error analysis to find out what domains, models, attack strategies, and decoding strategies affected the model’s performance and to what extent.
+This can be seen in Figure 2, Figure 3, Figure 4, and Figure 5.
+The texts were classified as machine generated if at least one percent of the tokens within the model’s context length were classified as machine generated.
+The F1 score obtained with the initial model trained on a single generator was 0.63, and the F1 score grew to 0.79 upon being trained on our full dataset.
+Evaluation was done directly without performing any preprocessing of the texts, and neither were our models trained on texts with any of the adversarial methods.
 Figure 2: F1 scores VS text sampling method used : Sampling strategy did not effect detection capability
-Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect.
+Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect than the corresponding base models.
 Figure 4: F1 scores VS each domain’s texts : news i.e the only training data domain was easier to detect.
 Figure 5: F1 scores VS adversarial method used in the input texts : homo-glyphs are the only real issue, and paraphrases to a small extent, while the rest can be handled through pre-processing.
-5.5 Comparisons with proprietary systems
+5.5 Comparisons with other systems
 While many proprietary systems claim to have excellent results, they often struggle with unseen domains and generators.
-Further many systems like ZeroGPT [^5] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^6] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API.
+Further, many systems like ZeroGPT [^6] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^7] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API.
 This would require manually evaluating the results through the UI by counting correct and incorrectly classified word counts.
-Hence, these systems were not used for comparisons for fine-grained scores.
-Several users have tried this manfully over a subset (Kadiyala, 2024) only to find a large gap in performance.
-For binary classifications as well, they have a minimum length threshold to even work, which would become incompatible over a large number of samples of the benchmark.
+Several users have tried this manually over a subset (Kadiyala, 2024) only to find a large gap in performance.
+However, for providing a comparison, we use other open-source systems for binary classification over the same benchmarks.
 6 Other Observations
-The sentences inside which text authorship switches from human to LLM or vice versa were found to be relatively shorter that the original text portions which they replaced.
+The sentences inside which text authorship switches from human to LLM or vice versa were found to be relatively shorter than the original text portions that they replaced.
 LLMs may be likely to finish the current sentence earlier than usual to move on to the next sentence in text completion scenarios.
-The mean length of the original portion and the replaced portions of those sentences for each language and generator can be seen in Table 5 and Table 6 respectively.
+The mean length of the original portion and the replaced portions of those sentences for each language and generator can be seen in Table 6 and Table 7, respectively.
 This observation was consistent across all languages and generators with a 20-30% reduction and a larger reduction in Hindi.
@@
 Table 8: Word level accuracy (.2f) of our models over our dataset (English)
-* excluding Chinese and Japanese
 7 Conclusion
-Despite not being trained on the domains or generators, the models built through our approach performed well over several benchmarks as seen in subsection 5.3 and subsection 5.4 over inputs which were from non-native speakers and consist of adversarial methods.
-Further, one case where many proprietary systems struggle is when the inputs were too short, which our models were able to overcome as seen in Figure 6, which demonstrates our models’ accuracy over our test set compared to input text’s sentence count.
-Table 7 displays our model’s performance over English subset of our dataset for each generator.
-A similar trend from subsection 5.4 was observed with models which are likely less instruction-tuned / not instruction-tuned tend to produce texts which are harder to distinguish than their alternatives.
+Despite not being trained on the domains or generators, the models built through our approach performed well over several benchmarks as seen in subsection 5.3 and subsection 5.4 over inputs that were from non-native speakers and consist of adversarial methods.
+Further, one case where many proprietary systems struggle is when the inputs were too short, which our models were able to overcome as seen in Figure 6, which demonstrates our models’ accuracy over our test set compared to the input text’s sentence count.
+Table 8 displays our model’s performance over the English subset of our dataset for each generator.
+A similar trend from subsection 5.4 was observed with models that are likely less instruction-tuned or not instruction-tuned; they tend to produce texts that are harder to distinguish than their alternatives.
 7.1 Scalability and scope for extension
-The original dataset used to train our current models as mentioned in section 3 consists of samples over 60 languages which would cover 70% of the world population’s primary language, and all of the languages are supported by existing multilingual transformer models making the process of scaling the work to more languages easier.
+The original dataset used to train our current models, as mentioned in section 3, consists of samples in over 60 languages, which would cover 70% of the world population’s primary language, and all of the languages are supported by existing multilingual transformer models, making the process of scaling the work to more languages easier.
 Despite not being trained on the generators or domains’ texts, our models were able to perform well on several benchmarks.
-Even reaching a F1 score of 0.79 against adversarial inputs while they were neither trained over them nor pre-processed.
-Similarly, creation and usage of such large datasets of other domains along with ours might result in robust and better models.
-We couldn’t explore the relation between instruction tuning sample size of LLMs and detectability of their texts due to the proprietary nature of most of the generators we used, but a similar study using open-data models could uncover more insights.
+They even reached an F1 score of 0.79 against adversarial inputs while they were neither trained over them nor pre-processed.
+Similarly, the creation and usage of such large datasets of other domains along with ours might result in robust and better models.
+We couldn’t explore the relation between the instruction tuning sample size of LLMs and the detectability of their texts due to the proprietary nature of most of the generators we used, but a similar study using open-data models could uncover more insights.
 7.2 Scope for Improvement
-As seen in Figure 5, almost none of the adversarial methods affected the models built through our approach other than paraphrasing and homo-glyphs.
-However homo-glyphs can be pre-processed by mapping them to the actual character they were imitating in the text.
-This would require a large collection of homo-glyph to character mapping set to use for pre-processing.
-Further, paraphrased samples of various number of iterations being included in the training dataset might lead to further improvements.
-It is also worth exploring how detectable are texts in cases where multiple generators contribute a portion each in a human authored text.
-Other missing adversarial methods that are likely to be used in practical scenarios include usage of proprietary systems that ’humanize’ a given text in an attempt to evade detection.
-7.3 Ideal Usage
-The models were built primarily for a human-in-the-loop use cases where the model would try to flag most of the likely machine-generated portions while the flagged content can be validated either through an ensemble of models or a human and hence a tilt towards higher recall can be observed in the metrics as seen in Table 13.
+As seen in Figure 5, almost none of the adversarial methods affected the models built through our approach other than paraphrasing and homoglyphs.
+However, homoglyphs can be pre-processed by mapping them to the actual character they were imitating in the text.
+This would require a large collection of homo-glyph-to-character mapping sets to use for preprocessing.
+Further, paraphrased samples of various numbers of iterations being included in the training dataset might lead to further improvements.
+It is also worth exploring how detectable texts are in cases where multiple generators contribute a portion each in a human-authored text.
+Other missing adversarial methods that are likely to be used in practical scenarios include the usage of proprietary systems that ’humanize’ a given text in an attempt to evade detection.
 Limitations
@@
 Furthermore, the experiments carried out did not include cases of multiple LLMs co-authoring a portion each of the same text.
-Acknowledgements
-Several stages of this work were supported by grants from Traversaal.ai, Cohere Labs & Microsoft Azure.
+The models were built primarily for a human-in-the-loop use cases where the model would try to flag most of the likely machine-generated portions while the flagged content can be validated either through an ensemble of models or a human and hence a tilt towards higher recall can be observed in the metrics
 Appendix A Pre- and Post- Boundary Comparisons
@@
 The hyper-parameters used for training the models can be seen in Table 11.
-We built a separate model for each language, the training was done over H100 SXM over 10h each.
+We built a separate model for each language, the training was done over H100 SXM over 237 hours.
 Appendix D Other Metrics
 The metrics over each type of text for each language and LLM separately can be seen in Table 15 to Table 26 respectively..
-Appendix E License
-The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0888https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.
 Language $\downarrow$ Accuracy Precision Recall F1-score